OpenStudy (anonymous):

What is the solution of X^2+64=0? -5 8 +-8 no solution

OpenStudy (anonymous):

Simplifying x2 + -64 = 0 Reorder the terms: -64 + x2 = 0 Solving -64 + x2 = 0 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '64' to each side of the equation. -64 + 64 + x2 = 0 + 64 Combine like terms: -64 + 64 = 0 0 + x2 = 0 + 64 x2 = 0 + 64 Combine like terms: 0 + 64 = 64 x2 = 64 Simplifying x2 = 64 Take the square root of each side: x = {-8, 8}

OpenStudy (anonymous):

x=8

OpenStudy (anonymous):

x=-8 too it's both

OpenStudy (anonymous):

^ :D

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Thanks:)

OpenStudy (anonymous):

Np.

OpenStudy (anonymous):

np :D

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

\[x^2+64=0 \] has only solutions in \[ \mathbb{C}\] though

OpenStudy (anonymous):

no because you subtract 64 then it's x^2 = 64 and then you square root it all :) which is x=8 or x=-8 :D

OpenStudy (anonymous):

ur finding x not b or c

OpenStudy (anonymous):

yes, can you take the square root from a negative number?

OpenStudy (anonymous):

Take a look at this: http://www.wolframalpha.com/input/?i=x%5E2%2B64%3D0

OpenStudy (anonymous):

i just know the answer i'm not really good at other things :/ sorry

OpenStudy (anonymous):

Anyway, in Reals this one has no answer, so -> no solution

OpenStudy (anonymous):

ummm idk O.o i think it's x= -8;8

OpenStudy (anonymous):

Just check then, if x=8 is your solution, plug it into your equation and tell me if that gives you zero. \[x^2+64=0\] so you say x=8 \[(8)^2+64=64+64=128 \neq 0 \]

OpenStudy (anonymous):

Now check for your other solution, you say x=-8 is a solution, so you plug that one in too: \[(-8)^2+64=64+64=128\neq0\]